# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 1.1.1. Large Deviations Principle
Let $D(t)=D(\mu_{t}\|\nu)$, $I(t)=I(\mu_{t}\|\nu)$, and $N(t)=\exp(-2D(\mu_{t}\|\nu))$ respectively be the relative entropy, Fisher information, and entropy power along the (generalized) heat flow $\partial_{t}\rho=L\rho$ with $L=\Delta-\nabla V\cdot\nabla$ and $\mu_{t}=\rho_{t}\nu$. It is easy to see that $\Theta(\alpha)$ grows exponentially as $\alpha\to\infty$, as long as $N(t)$ is concave in the time $t$ and $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$. This is because, in this case, $D(t)$ continuously decays to zero as $t\to\infty$ and $N^{\prime}(t)=2N(t)I(t)$ is nonincreasing, which implies that for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, we always have $N(0)I(0)\geq N(t_{0})I(t_{0})\geq\exp(-2\alpha_{0})\Theta(\alpha_{0})$ with $t_{0}$ being the solution to $D(t_{0})=\alpha_{0}$. Therefore, $I(0)\geq\Theta(\alpha_{0})\exp(2(D(0)-\alpha_{0}))$ for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, i.e., exponential growth of $\Theta$: $\Theta(\alpha)\geq\Theta(\alpha_{0})\exp(2(\alpha-\alpha_{0}))$ for all $\alpha\geq\alpha_{0}$. The concavity of the entropy power holds under the $\mathrm{CD}(0,m)$-condition with $m\in[n,\infty)$, as shown by S. Li and X. Li [42, Theorem 2.2]. The condition that $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$ is guaranteed as long as a standard log-Sobolev inequality (or a Poincaré inequality) is admitted. These two conditions are satisfied by the (normalized) unweighted measure on a bounded manifold with nonnegative Ricci curvature, which hence admits exponential growth of $\Theta$ and also exponential growth of $\Lambda$ (since Assumptions 1 and 2 are satisfied in this setting and Theorem 1 is valid). See other sufficient conditions to ensure exponential growth of $\Theta$, e.g. , in [7, Theorem 6.8.1][42, Theorem 2.5].
@@
Consider the hypercube $(0,1)^{n}$ with Euclidean metric and uniform measure. The standard log-Sobolev inequality and the Poincaré inequality with constants $K_{LS}=K_{P}=\pi^{2}$ ($K_{LS},K_{P}$ defined later in (1.12) and (1.18)) respectively hold in this setting [27], which implies $\lim_{\alpha\to 0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}=\pi^{2}$. Moreover, the $\mathrm{CD}(0,m)$-condition holds obviously. Thus, as mentioned in Example 1, we have exponential growth of $\Theta$, i.e., $\Theta(\alpha)\geq 2\pi^{2}\alpha_{0}\cdot e^{2(\alpha-\alpha_{0})}\geq\frac{\pi^{2}}{e}\cdot e^{2\alpha}$ (with $\alpha_{0}$ chosen $1/2$), which further implies exponential growth of $\Lambda$, i.e., $\Lambda(\alpha)\geq\sqrt{\frac{\pi^{2}}{e}}\cdot e^{\alpha}$. This lower bound is sharp in the exponent by verifying balls centered at the vertices of the hypercube, since for these balls with volume $e^{-n\alpha}$, their surface areas are $(1+o(1))\sqrt{\frac{\pi en}{2}}\cdot e^{\alpha(1-n)}$ as $n\to\infty$, giving $\Lambda(\alpha)\leq\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$. In fact, it was already known [58, 29] that given each $n$, for all small volumes, balls centered at the vertices are isoperimetric minimizers. However, it is unclear whether “small volume” here is large enough to include “exponentially small volume”. We conjecture the following precise relation holds: $\Lambda(\alpha)\sim\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$, or equivalently, $\breve{\Theta}(\alpha)\sim\frac{\pi e}{2}\cdot e^{2\alpha}$ as $\alpha\to\infty$, suggesting that half-normal distributions truncated on $(0,1)$ with sufficiently small variances approximately saturate nonlinear log-Sobolev inequalities for large $\alpha$. The factor $\frac{\pi^{2}}{e}$ obtained above is about $0.85\cdot\frac{\pi e}{2}$.
@@
Showing the equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature on this topic are known only in a weak sense. In particular, it was proved by Ledoux [36, 37] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [14] on Euclidean spaces with any log-concave probability measures that for a fixed $n$,
$I_{n}(a)\geq a\sqrt{2K\ln\frac{1}{a}}$
holds for some $K>0$, if and only if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense for product spaces, which showing a precise equivalence between the isoperimetric inequality and (nonlinear) log-Sobolev inequality. Our equivalence will be applied in the following to compare the optimal constants in various inequalities.
@@ 1.1.2. Companion of Various Constants
While Theorem 1 is primarily tailored to subsets with probability measures decaying exponentially in $n$, it admits application to bounding the isoperimetric profile for subsets with probability measures decaying arbitrarily slowly. In other words, Theorem 1 can be used to quantitatively compare the log-Sobolev constant, the isoperimetric constant, and other constants for various inequalities. To this end, we define the following constants.
  • $K_{CD}$ denotes the optimal constant $K$ in the Ricci curvature-dimension condition
@@
    $K_{LS}^{+}:=\sup_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to\infty}\frac{\breve{\Theta}(\alpha)}{2\alpha}.$
@@
    $\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\quad\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},$
@@
    $\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},\quad\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})\textrm{ s.t. }D(\mu\|\nu)\geq D_{0}.$
@@
Under Assumption 1 (i.e., $K_{CD}\geq 0$), the following hold.
  1. (1)
    It holds that
@@
    and
    $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{T}^{+},$
    where $x\wedge y=\min\{x,y\}$. Under the additional Assumption 2, the equality in (d) holds and moreover,
    $K_{LS}^{+}\overset{\mathrm{(l)}}{\leq}K_{IS}^{+}.$
    In addition, if $K_{CD}$ is equal to any one of $K_{IS}^{-},K_{IS},K_{LS},K_{HC},K_{T},K_{C}$, then all the inequalities in (a)–(g) of (1.14) become equalities.
  2. (2)
    For any log-concave probability measure $\nu$ on $\mathbb{R}$ with density $e^{-V}$ such that $\lim_{x\to\infty}V^{\prime\prime}(x)$ and $\lim_{x\to-\infty}V^{\prime\prime}(x)$ exist,
    $K_{IS,1}^{+}:=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq K_{LS}^{+}.$
For Statement 1, Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [33]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [54]. Equality (g) is also well known; see e.g. [65, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). Inequality (k) follows similarly to $K_{LS}\leq K_{T}$. Inequality (l) follows by Theorem 1, with a detailed proof given in Appendix A. So, it only remains to prove Inequality (b). This is done in Appendix A. In addition, a weaker version of (b), $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$, was already known; see Theorem 22.21 of [65]. However, we do not know if this is equivalent to the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds. Statement 2 is proven in Appendix A.
@@
where the density $1+\epsilon\cos\pi x$ asymptotically attains the constants $K_{LS}$ and $K_{T}$ as $\epsilon\to 0$. Moreover, it holds that $\sqrt{2\pi}\leq K_{IS}^{-}\leq\pi^{2}$; the lower bound will be explaind in Example 4. However, the exact value of $K_{IS}^{-}$ is still unknown.
Without Assumption 2, we have the following weaker result, as a consequence of Theorem 2.
@@
Under Assumption 1, it holds that for any $\alpha>0$,
(1.15) $\sqrt{\frac{K_{LS}}{4}\alpha}\leq\sqrt{\frac{K_{T}}{4}\alpha}\leq\underline{\Lambda}(\alpha)\leq\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$}.
1.1.3. Central Limit Regime
Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(x,y)=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces. The proof is provided in Appendix B.
Theorem 4.
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
@@ 1.1.3. Central Limit Regime
where $I_{\mathrm{G}}$ is the Gaussian isoperimetric profile and $K_{0}:=\inf_{a\in(0,1)}\frac{I_{\nu}(a)}{I_{\mathrm{G}}(a)}$ [13, 8]. An upper bound can be obtained by considering the $\mathbf{f}$-coordinate half-space given in (1.8), as shown in the following theorem. The proof is based on the coarea formula and the central limit theorem, which is given in Appendix C.
@@
Assume that $I_{\inf}$ is continuous (e.g., Assumption 1 holds). Then, for any $a\in[0,1]$,
@@
Thus, $K_{LS}=K_{P}$ if $\Theta(\alpha)\geq 2K_{P}\alpha$ (e.g., if $\Theta$ is convex), and otherwise, $K_{LS}<K_{P}$ (e.g., for exponential measures [65, Example 21.19]). For the latter case, the upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ is not tight for sufficiently small $a$. Combining the best known lower bound in (1.16), the upper bound in (1.17), and the trivial upper bound $I_{1}$ yields
@@
1.2. Open Problems
As mentioned above, one major open question on the topic considered in this paper is the determination of $K_{IS}^{-}$.
Problem 1.
Under Assumption 1, is it true that $K_{IS}^{-}=K_{LS}$?
A more challenging problem is to exactly determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
@@ 1.2. Open Problems
Under Assumption 1, what is the explicit expression of $I_{\inf}$?
Another interesting topic is to remove Assumption 2 from Theorem 1. To this end, we propose the following conjecture.
Conjecture 1 (Pythagorean-Like Inequality).
Let $\nu_{1}=e^{-V_{1}}\mathrm{vol}_{1},\nu_{2}=e^{-V_{2}}\mathrm{vol}_{2}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Then,
(1.19) $\left(\frac{I_{\nu_{1}\otimes\nu_{2}}(a)}{a}\right)^{2}\geq\inf_{a_{1}a_{2}=a}\left(\frac{I_{\nu_{1}}(a_{1})}{a_{1}}\right)^{2}+\left(\frac{I_{\nu_{2}}(a_{2})}{a_{2}}\right)^{2},\;\forall a\in(0,1).$
Remark 5.
A weaker version of this conjecture was proven by Grigor’yan [32] and Morgan [50].
This conjecture can be equivalently formulated as the plane version, as explained in the following. Any probability measure $\nu$ satisfying Assumption 1 admits a concave isoperimetric profile $I_{\nu}$. For every such probability measure $\nu$, one can find a unique symmetric (w.r.t. $x=0$) log-concave probability measure $\tilde{\nu}$ on $\mathbb{R}$ such that $I_{\tilde{\nu}}=I_{\nu}$ [14]. Explicitly, the cdf $F$ of $\tilde{\nu}$ is exactly the inverse of the function
$G(y)=\int_{1/2}^{y}\frac{dz}{I_{\nu}(z)},$
which will induces $I_{\tilde{\nu}}=f\circ F^{-1}=I_{\nu}$, with $f=F^{\prime}$ denoting the pdf of $\nu$. Let $\nu_{1}=e^{-V_{1}}\mathrm{vol},\nu_{2}=e^{-V_{2}}\mathrm{vol}$ with $V_{1}\in C^{2}(\mathcal{M}_{1}),V_{2}\in C^{2}(\mathcal{M}_{2})$ be probability measures respectively on Riemannian manifolds $\mathcal{M}_{1},\mathcal{M}_{2}$. Assume both $V_{1}$ and $V_{2}$ satisfy Assumption 1. Let $\tilde{\nu}_{1},\tilde{\nu}_{2}$ be the symmetric log-concave probability measures on $\mathbb{R}$ induced by $\nu_{1},\nu_{2}$ respectively such that $I_{\tilde{\nu}_{i}}=I_{\nu_{i}},i=1,2$. Denote $F_{i},f_{i}$ as the cdf and pdf of $\tilde{\nu}_{i}$, $i=1,2$. Define
(1.20) $J_{1}(a)$ $=\inf_{\sigma}\int_{\mathcal{M}_{1}}\sqrt{I_{\nu_{2}}^{2}(\sigma(x))+|\nabla\sigma(x)|^{2}}{\text{\rm d}}\nu_{1}(x),$
$J_{2}(a)$ $=\inf_{\varphi}\int_{0}^{1}\sqrt{I_{\nu_{2}}^{2}(\varphi(t))+I_{\nu_{1}}^{2}(t)\varphi^{\prime}(t)^{2}}{\text{\rm d}}t,$
$J_{3}(a)$ $=\inf_{g}\int_{\mathbb{R}}\sqrt{1+(g^{\prime}(x))^{2}}f_{2}(g(x))f_{1}(x){\text{\rm d}}x,$
where the first infimum is taken over all smooth functions $\sigma:\mathcal{M}_{1}\to[0,1]$ such that $\int_{\mathcal{M}_{1}}\sigma(x){\text{\rm d}}\nu_{1}(x)=a$, the second infimum is taken over all smooth functions $\varphi:[0,1]\to[0,1]$ such that $\int_{0}^{1}\varphi(t){\text{\rm d}}t=a$, and the third infimum is taken over all smooth functions $g:\mathbb{R}\to\mathbb{R}$ such that $\int_{\mathbb{R}}F_{2}(g(x))f_{1}(x){\text{\rm d}}x=a$. The following proposition shows that the isoperimetric profile of $\nu_{1}\otimes\nu_{2}$ will decrease if we replace one or both of $\nu_{1},\nu_{2}$ to the corresponding $\tilde{\nu}_{1},\tilde{\nu}_{2}$. This proposition is an extension of the Gaussian version due to Bobkov [13] and the weighted Riemannian slab version due to [46] to Riemannian manifolds, and a simple proof is provided in Appendix D for completeness.
Proposition 2.
It holds that for all $a\in[0,1]$,
$I_{\nu_{1}\otimes\nu_{2}}(a)$ $\geq I_{\nu_{1}\otimes\tilde{\nu}_{2}}(a)=J_{1}(a)\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)=J_{2}(a)=J_{3}(a).$
Remark 6.
By symmetry, $I_{\nu_{1}\otimes\nu_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\nu_{2}}(a)\geq I_{\tilde{\nu}_{1}\otimes\tilde{\nu}_{2}}(a)$ also holds. In addition, this proposition can be used to derive dimension-dependent isoperimetric inequalities. Suppose that $\Upsilon_{i}:[0,1]\to[0,\infty),i\in\{1,2,...,n\}$ are functions such that $I_{\nu_{1}}\geq\Upsilon_{1}$ and for all $i\in\{2,3,...,n\}$,
$\int_{\mathcal{M}_{i}}\sqrt{\Upsilon_{i-1}^{2}(\sigma_{i}(x))+|\nabla\sigma_{i}(x)|^{2}}{\text{\rm d}}\nu_{i}(x)\geq\Upsilon_{i}(\int_{\mathcal{M}_{i}}\sigma_{i}(x){\text{\rm d}}\nu_{i}(x))$
holds for all smooth functions $\sigma_{i}:\mathcal{M}_{i}\to[0,1]$. Then, by repeatly applying Proposition 2, $I_{\nu_{1}\otimes\nu_{2}\otimes...\otimes\nu_{i}}\geq\Upsilon_{i},i\in\{1,2,...,n\}$.
This proposition implies that it suffices to restrict the inequality in Conjecture 1 to the Euclidean plane, yielding the following equivalent formulation of Conjecture 1.
Conjecture 2.
For any symmetric log-concave probability measures $\nu_{1},\nu_{2}$ on $\mathbb{R}$ (in which case, $I_{\nu_{i}}=f_{i}\circ F_{i}^{-1}$ with $f_{i}$ and $F_{i}$ denoting the pdf and cdf of $\nu_{i}$),
$\left(\frac{I_{\nu_{1}\otimes\nu_{2}}(a)}{a}\right)^{2}\geq\inf_{a_{1}a_{2}=a}\left(\frac{I_{\nu_{1}}(a_{1})}{a_{1}}\right)^{2}+\left(\frac{I_{\nu_{2}}(a_{2})}{a_{2}}\right)^{2},\;\forall a\in(0,1).$
We next show that under Conjecture 2, Theorem 1 still holds for one-dimensional Euclidean space without Assumption 2. The proof of the following proposition is given in Appendix E.
Proposition 3.
Assume that Conjecture 2 holds. Let $\nu$ be a log-concave probability measure on $\mathbb{R}$ with density $e^{-V}$ such that $\lim_{x\to\infty}V^{\prime\prime}(x)$ and $\lim_{x\to-\infty}V^{\prime\prime}(x)$ exist, Theorem 1 still holds.

### v4

@@ 1.1.1. Large Deviations Principle
Let $D(t)=D(\mu_{t}\|\nu)$, $I(t)=I(\mu_{t}\|\nu)$, and $N(t)=\exp(-2D(\mu_{t}\|\nu))$ respectively be the relative entropy, Fisher information, and entropy power along the (generalized) heat flow $\partial_{t}\rho=L\rho$ with $L=\Delta-\nabla V\cdot\nabla$ and $\mu_{t}=\rho_{t}\nu$. It is easy to see that $\Theta(\alpha)$ grows exponentially as $\alpha\to\infty$, as long as $N(t)$ is concave in the time $t$ and $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$. This is because, in this case, $D(t)$ continuously decays to zero as $t\to\infty$ and $N^{\prime}(t)=2N(t)I(t)$ is nonincreasing, which implies that for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, we always have $N(0)I(0)\geq N(t_{0})I(t_{0})\geq\Theta(\alpha_{0})e^{-2\alpha_{0}}$ with $t_{0}$ being the solution to $D(t_{0})=\alpha_{0}$. Therefore, $I(0)\geq\Theta(\alpha_{0})e^{2(D(0)-\alpha_{0})}$ for any $\mu_{0}$ satisfying $D(0)\geq\alpha_{0}$, i.e., exponential growth of $\Theta$ (when $\Theta(\alpha_{0})>0$): $\Theta(\alpha)\geq\Theta(\alpha_{0})e^{2(\alpha-\alpha_{0})}$ for all $\alpha\geq\alpha_{0}$. The concavity of the entropy power holds under the $\mathrm{CD}(0,m)$-condition with $m\in[n,\infty)$ for smooth Riemannian manifolds, as shown by S. Li and X. Li [47, Theorem 2.2], and also under the $\mathrm{RCD}(0,m)$-condition with $m\in[n,\infty)$ and $\int_{\mathcal{X}}\left[\frac{|\Delta\rho|^{2}}{\rho}+\frac{|\nabla\Delta\rho|^{2}}{\rho}\right]{\text{\rm d}}\nu<\infty$ for Polish metric measure spaces with $\rho$ denoting the heat flow, as shown by X. Li and E. Zhang [48, Theorem 3.1 and Remark 3.3]. The condition that $\Theta(\alpha_{0})>0$ for some $\alpha_{0}>0$ is guaranteed as long as a standard log-Sobolev inequality (or a Poincaré inequality) is admitted. These two conditions, concavity of $N$ and positivity of $\Theta(\alpha_{0})$, are satisfied by the (normalized) unweighted measure on a bounded manifold with nonnegative Ricci curvature, which hence admits exponential growth of $\Theta$ and also exponential growth of $\Lambda$ (since Conditions 1 and 3 are satisfied in this setting and Theorem 1 is valid). See other sufficient conditions to ensure exponential growth of $\Theta$ in [10, Theorem 6.8.1][47, Theorem 2.5].
@@
Consider the hypercube $0,1]^{n}$ (equivalently, $(0,1)^{n}$) with Euclidean metric and uniform measure. The standard log-Sobolev inequality and the Poincaré inequality with constants $K_{LS}=K_{P}=\pi^{2}$ ($K_{LS},K_{P}$ defined later in (1.14) and (1.20)) respectively hold in this setting [32], which implies $\lim_{\alpha\to 0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to 0}\frac{\Theta(\alpha)}{2\alpha}=\pi^{2}$. Moreover, the $\mathrm{RCD}(0,m)$-condition holds obviously for $m>n$ and the condition $\int_{[0,1]}\left[\frac{|\Delta\rho|^{2}}{\rho}+\frac{|\nabla\Delta\rho|^{2}}{\rho}\right]{\text{\rm d}}\nu<\infty$ is satisfied for any $t>0$ for heat flow $\rho$ with the Neumann boundary condition [67, Section 4.2] (since $[0,1]$ is compact, $\rho\in C^{\infty}([0,1])$, and $\rho_{t}>0$ for any $t>0$). Thus, as mentioned in Example 2, we have exponential growth of $\Theta$, i.e., $\Theta(\alpha)\geq 2\pi^{2}\alpha_{0}\cdot e^{2(\alpha-\alpha_{0})}\geq\frac{\pi^{2}}{e}\cdot e^{2\alpha}$ (with $\alpha_{0}$ chosen as $1/2$), which further implies exponential growth of $\Lambda$, i.e., $\Lambda(\alpha)\geq\sqrt{\frac{\pi^{2}}{e}}\cdot e^{\alpha}$. This lower bound is sharp in the exponent by verifying balls centered at the vertices of the hypercube, since for these balls with volume $e^{-n\alpha}$, their surface areas are $(1+o(1))\sqrt{\frac{\pi en}{2}}\cdot e^{\alpha(1-n)}$ as $n\to\infty$, giving $\Lambda(\alpha)\leq\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$. In fact, it was already known [63, 35] that given each $n$, for all small volumes, balls centered at the vertices are isoperimetric minimizers. However, it is unclear whether “small volume” here is large enough to include “exponentially small volume”. We conjecture the following precise relation holds: $\Lambda(\alpha)\sim\sqrt{\frac{\pi e}{2}}\cdot e^{\alpha}$, or equivalently, $\breve{\Theta}(\alpha)\sim\frac{\pi e}{2}\cdot e^{2\alpha}$ as $\alpha\to\infty$, suggesting that half-normal distributions truncated on $(0,1)$ with sufficiently small variances approximately saturate nonlinear log-Sobolev inequalities for large $\alpha$. The factor $\frac{\pi^{2}}{e}$ obtained above is about $0.85\cdot\frac{\pi e}{2}$.
@@
Showing equivalence among inequalities from various fields is an interesting topic, since it will unveil the intrinsic relationships among diverse branches of mathematics. However, most of equivalences in the literature on this topic are known only in a weak sense. In particular, it was proved by Ledoux [41, 42] on Gaussian spaces or compact Riemannian manifolds and by Bobkov [20] on Euclidean spaces with log-concave probability measures that for a fixed $n$, a Gaussian-type isoperimetric inequality
$I_{n}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a$
holds for some $K>0$, if and only if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$. Theorem 1 provides an equivalence in a stronger sense, which showing a precise equivalence between the isoperimetric inequality and (nonlinear) log-Sobolev inequality. As demonstrated in the case of bounded manifolds, Ledoux and Bobkov’s weak equivalence exhibits significant looseness for small $a$ (that is, in the context of small sets).
Although our equivalence is proven for product spaces, it in fact implies Ledoux and Bobkov’s weak equivalence for non-product spaces. In particular, by the facts that $I_{1}\geq I_{n}$ and $\mathcal{M}$ is an arbitrary (non-product) space, if the standard log-Sobolev inequality holds with some constant $K^{\prime}>0$, then our equivalence implies that $\liminf_{a\to 0}\frac{I_{1}(a)}{I_{\mathrm{G}}(a)}\geq\liminf_{n\to\infty}\frac{I_{n}(a_{n})}{I_{\mathrm{G}}(a_{n})}=\sqrt{K^{\prime}},$ where $a_{n}=e^{-n\alpha}$ for any $\alpha>0$. (More details on the inequality here is given in the following subsubsection.) That is, $I_{1}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a$ for some $K>0$, yielding a Gaussian-type isoperimetric inequality for the space $\mathcal{M}$. Conversely, if a Gaussian-type isoperimetric inequality for $\mathcal{M}$ holds with constant $K>0$, then by the well known fact that $I_{1}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a\Longleftrightarrow I_{n}(a)\geq\sqrt{K}I_{\mathrm{G}}(a),\forall a,n$, our equivalence implies the standard log-Sobolev inequality with the same constant $K$.
Our equivalence will be applied in the following to compare the optimal constants in various inequalities.
@@ 1.1.2. Companion of Various Constants
While Theorem 1 is primarily tailored to subsets of weighted volumes decaying exponentially in $n$, it admits application to bounding the isoperimetric profile for weighted volumes decaying arbitrarily slowly or even constant weighted volumes. In other words, Theorem 1 can be used to quantitatively compare the log-Sobolev constant, the isoperimetric constant, and other constants for various inequalities. To this end, we define the following constants.
  • $K_{CD}$ denotes the optimal constant $K$ in the Bakry–Émery–Ricci curvature-dimension condition
@@
    (1.15) $K_{LS}^{+}:=\sup_{\alpha>0}\frac{\breve{\Theta}(\alpha)}{2\alpha}=\lim_{\alpha\to\infty}\frac{\breve{\Theta}(\alpha)}{2\alpha}.$
@@
    $\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),\quad\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},$
@@
    $\mathsf{W}_{2}(\mu,\nu)\leq\sqrt{\frac{2D(\mu\|\nu)}{K}},\quad\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})\textrm{ s.t. }D(\mu\|\nu)\geq D_{0}.$
@@
Suppose the weighted Riemannian manifold $(\mathcal{M},d,\nu)$ satisfies Condition 1 (i.e., $K_{CD}\geq 0$). Then, it holds that
@@
and $K_{LS}\overset{\mathrm{(j)}}{\leq}K_{LS}^{+}.$ If additionally, Condition 3 holds when the condition in (1.10) holds, then the equality in (d) holds and moreover, $K_{LS}^{+}\overset{\mathrm{(k)}}{\leq}K_{IS}^{+}.$ In addition, if $K_{CD}$ is equal to any one of $K_{IS}^{-},K_{IS},K_{LS},K_{HC},K_{T},K_{C}$, then all the inequalities in (a)–(g) of (1.17) become equalities.
Inequality (a) is trivial, by noting that $0\leq K_{CD}/K_{C}\leq 1$. Inequality (c) is due to the fact that $n\mapsto I_{n}(a)$ is nonincreasing. Equality (d) is implied by Theorem 1. Equality (e) is a classic result of Gross [38]. The inequality $K_{LS}\leq K_{T}$ (or equivalently, Inequality (f)) was proven by Otto and Villani [58]. Equality (g) is also well known; see e.g. [73, Remark 22.23]. Inequalities (h) and (j) follow by definition. Equality (i) follows similarly to (g). Inequality (k) follows by Theorem 1, with a detailed proof given in Appendix C. So, it only remains to prove Inequality (b). This is done in Appendix C. In addition, a weaker version of (b), $\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}\leq K_{LS}$, was already known; see Theorem 22.21 of [73]. However, we do not know if this is equivalent to the inequality (b) since we do not know if $K_{IS}^{-}=K_{IS}$ holds.
@@
where the density $1+\epsilon\cos\pi x$ asymptotically attains the constants $K_{LS}$ and $K_{T}$ as $\epsilon\to 0$. Moreover, it holds that $\sqrt{2\pi}\leq K_{IS}^{-}\leq\pi^{2}$; the lower bound will be explained in Example 5. However, the exact value of $K_{IS}^{-}$ is still unknown.
An open problem on this topic is the following one.
Problem 1.
Under Condition 1, is it true that $K_{IS}^{-}=K_{LS}$?
1.1.3. Central Limit Regime
We now consider the asymptotics of the isoperimetric profile for fixed weighted volume as $n\to\infty$. Let $I_{\infty}$ be the isoperimetric profile defined for the infinite dimensional product space $\mathcal{M}^{\infty}$ equipped with the product topology and the product measure $\nu_{\infty}:=\nu^{\otimes\infty}$, but the enlargement of a set is defined under $d_{\infty}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}}$ (although it might diverse to infinity for two arbitrary points). Then, it is obvious that $I_{\inf}(a)\geq I_{\infty}(a)$. In the following, we show that this is actually an equality. That is, the isoperimetric profile $I_{\infty}$ on the infinite dimensional space is the pointwise limit of the isoperimetric profiles on finite dimensional spaces. The proof is provided in Appendix D.
@@ 1.1.3. Central Limit Regime
Assume that $I_{\inf}$ is continuous (e.g., Condition 1 holds). Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
@@
where $I_{\mathrm{G}}$ is the Gaussian isoperimetric profile and $K_{0}:=\inf_{a\in(0,1)}\frac{I_{\nu}(a)}{I_{\mathrm{G}}(a)}$ [19, 11]. An upper bound can be obtained by considering the $\mathbf{f}$-coordinate half-space given in (1.8), as shown in the following theorem. The proof is based on the coarea formula and the central limit theorem, which is given in Appendix E.
@@
Assume that $I_{\inf}$ is continuous (e.g., Condition 1 holds). Then, for any $a\in[0,1]$,
@@
Thus, $K_{LS}=K_{P}$ if $\Theta(\alpha)\geq 2K_{P}\alpha$ (e.g., if $\Theta$ is convex), and otherwise, $K_{LS}<K_{P}$ (e.g., for exponential measures [73, Example 21.19]). For the latter case, the upper bound $\sqrt{K_{P}}I_{\mathrm{G}}(a)$ is not tight for sufficiently small $a$. Combining the best known lower bound in (1.18), the upper bound in (1.19), and the trivial upper bound $I_{1}$ yields
@@
A challenging problem is to exactly determine the function $I_{\inf}$ (or equivalently, $I_{\infty}$). This problem is a commonly recognized open one.
@@
Under Condition 1, what is the explicit expression of $I_{\inf}$?